# Rigor in AI: Doing Rigorous AI Work Requires a Broader, Responsible AI-Informed Conception of Rigor

### v1

@@ 2.2 Normative rigor
Because of differences in underlying evidentiary standards, background assumptions, goals and ideals, and theoretical frameworks, different disciplines and communities have different ways of determining what methods to use, what types of evidence are needed or sufficient, and why the resulting work or resulting knowledge matters [29, 34]. The assumptions and theories that underpin our work are not derived “out of thin air, but as a function of our experiences with the world in general and specifically with our colleagues, through dialogue, and the literature” [58], all of which can shape the type, direction, and quality of research [67, 42, 54]. Disciplinary and community norms are intertwined with a researcher’s personal norms and drivers [43], what catches their interest or advances their goals. Making explicit how this mix of influences shapes the AI community’s work can help others understand them and enable debates on their appropriateness.
@@
Consider the emerging research on developing AI personas as tools to simulate users and human study participants [e.g., 6, 82, 51, 108]. While this research reflects common norms and beliefs in AI communities around scalability and efficiency [e.g., 20, 3], by aiming to replace human participants these tools may fail to align or may even conflict with foundational values and norms around representation, participation, inclusion, and understanding [3, 146] that underpin many types of work this research seeks to support, such as user experience or social science studies with human participants. Such value conflicts “cannot be alleviated with better training or improved model performance alone” [3].
@@
Common beliefs and expectations in the AI community—such as the goal of producing generalizable findings that are often abstracted away from any use context [20, 68]—have led to a reduction of problems and use scenarios “to a common set of representations or affordances” [20], and thus (even if implicitly) to de-valuing how context shapes datasets and tools. Reliance on undisclosed normative assumptions about problem statements, artifacts, or constructs that are contested or value-laden, yet treated as if they are neutral or generally applicable, can however be particularly misleading when a “field employs the language of procedural adherence to project a sense of certainty, objectivity, and stability” [56]. While dominant values in AI like performance, generalization, and efficiency [20] may encourage extensive evaluations across as many metrics, benchmarks, and baselines as possible [155, 20], these evaluations also tell us little about when, whether, or which improvements are necessary, desirable, or translate to meaningful benefits when deployed. Such expectations and norms—along resource-related considerations and constraints such as about costs and strict timeliness [153, 155, 61]—nonetheless influence what type of work gets prioritized and done [48].
@@
Mechanisms to promote normative rigor All research is shaped by normative considerations. Normative rigor asks us to make these considerations explicit—such as via ethical and positionality statements [107] or other disclosure practices—and can include aspects related to expectations around the significance and impact of research [e.g., 104, 58] or what ethical norms to follow [e.g., 106].
@@
Research significance:  Central to normative considerations are expectations about how our work and its outcomes should intervene in the world, or why the work matters and is appropriate. The goal of research is often to “generate knowledge that will have positive practical impact” [10] or that advances the field [104], with norms around what work is worthwhile driven by desires to reward such work [148, 32]. There is however also an increasing recognition that the appraisal of the work’s quality and significance should also include the work’s potential for harm and not only for benefits [58, 27, 10]. Such appraisal typically rests on claims about the possible impacts from either the work’s process or its outcomes. This echoes the concept of consequential validity from social science measurement scholarship [97, 142], which asks us when determining a measurement instrument’s (in)validity to consider the consequential basis both of its use and of possible inferences (along with actions those inferences may entail) [97, 70]. Reckoning with the impact of our work thus promotes not only discussions about whether current norms around what constitutes good work are appropriate, but also promotes methodological (§2.4) and interpretative rigor (§2.6).
@@
Positionality and ethical statements:  Researchers’ personal, disciplinary and institutional backgrounds, their lived experiences, and their goals motivate and shape how they approach their work. Positionality statements are a mechanism meant to make such considerations explicit in order to help others contextualize the research and research outcomes [107, 83]. Positionality, however, does not only encompass aspects related to researchers’ beliefs and values, but also those related to what knowledge they draw on, how they know what they know, and how they make methodological choices [17]. Thus, it can also aid or compromise epistemic (§2.1) and methodological rigor (§2.4). Ethical statements can further complement positionality statements by foregrounding the ethical concerns researchers grappled with or mitigated before or while conducting the work [107]. Yet this practice of disclosing whether such concerns were considered and how they shaped any methodological choices (if at all) remains inconsistently adopted across AI communities.
@@ 2.3 Conceptual rigor
Assume a researcher wishes to evaluate whether a model “hallucinates.” In AI research, the construct of “hallucination” has, however, been used to refer to several distinct types of system behaviors [93], including cases of generating content which is nonsensical, which contains factual errors, which is not in the input data, which is not in the training data, or a mix of these behaviors. Further, all these different understandings of what it means for a model to “hallucinate” are markedly different from the more common use of the term that requires an ability to perceive, feel, or have sensory experiences [e.g., 38], and the term can thus carry meanings incompatible with AI systems. To understand what a researcher’s evaluation of whether the model “hallucinates” means, we need to know which conceptualization they use, and if that conceptualization is sensible.
@@
While contested constructs—those that have competing or even conflicting definitions, like “hallucination,” “value alignment,” “AGI,” or “human-likeness”—are increasingly common in AI research and practice, their definitions often remain elusive, ambiguous, or poorly specified [e.g. 26, 142, 23]. Without clarity about what specifically we are analyzing, measuring, or striving for, it can be hard to assess progress or make any useful or reliable claims. Work can also rely on constructs that are inconsistent with any theoretical tradition, such as treating identity categories as fixed and objective rather than continuous and constructed [90] or more generally failing to recognize how some constructs are innately fluid, non-deterministic, and fuzzy [19]. Such a lack of clarity can also hinder replicability and reproducibility or facilitate speculative post-factum interpretations, yielding possibly unsound and unfounded claims, or a conflation of proxy measurements (which may or may not measure any version of the underlying construct) with the construct under analysis—thus also undermining epistemic (§2.1), methodological (§2.4), and interpretative rigor (§2.6).
@@
Mechanisms to promote conceptual rigor Conceptual rigor requires attention to conceptual clarity—which construct we are after and how it is defined, appropriate conceptual systematization—the process by which the definition is made specific, and terminological rigor—that the terms used to refer to a construct do not harbor meanings that can lead to a misinterpretation of what the construct is.
@@
Conceptual clarity:  A growing number of objects in AI research are ambiguous or poorly specified, or else are objects for which we lack consensus about what they are or what they are for. The examination by Saphra and Wiegreffe [122] of what is meant by “mechanistic interpretability” is instructive: not only does the term have multiple competing meanings, but those meanings also reflect distinct disciplinary orientations and epistemic origins; a lack of clarity about which meaning is under use can obfuscate not only what the work does, but also why and how the work is done. Similar critiques have been made about the lack of conceptual clarity about what unlearning [40], bias [25], model collapse [125], interpretability [86], or generalization [87] mean. While we see growing concerns about the lack of conceptual clarity surrounding many aspects of AI research, from how desired capabilities are described or what metrics to optimized for [71, 26, 62, 142, 122], these remain largely overlooked in discussions about research integrity and quality in AI.
@@
Conceptual systematization:  In practice many constructs involve a “broad constellation of meanings and understandings” [1], and working with them requires making choices about which meanings to use, “narrowing [them] into an explicit definition” [142]. The process of conceptual systematization asks researchers to engage not only with a high-level construct in the abstract, but to grapple more concretely with what it means in the context of the work they are conducting and how it relates to empirical observations or other constructs. Conceptual systematization is a prerequisite for rigorous measurement, (computational) specification, empirical analysis, and theory development [e.g., 142, 110, 1, 62], and is thus a prerequisite for methodological rigor (§2.4).
@@
Terminological rigor:  Conceptual rigor depends on terminological choices and what those choices communicate. Many terms in AI often carry over meanings from the human realm or other disciplinary contexts that are incompatible with AI systems, and can mislead [36, 87, 24, 115], suggest “unproven connotations,” or lead to “collisions with other definitions, or conflation with other related but distinct concepts” [87]. Blili-Hamelin et al. [24] note that “when researchers equate human faculties with model proxies… [t]his rhetorical move is enabled by using colloquial terms like ‘imagination’ without considering whether it corresponds to the human faculty,” leading to inflated claims. Clear, precise language “help[s] dispel speculative, scientifically unsupported portrayals of [AI] systems, and support more factual descriptions of them” [37], and thus clear scientific communication.
@@ 2.4 Methodological rigor
Rigor is often “conceptualized as the appropriate execution of [methods]” [104], with discussions about rigor in AI centering around methodological considerations, from data collection and analysis, to model training and tuning, to experimental practices [e.g., 62, 129, 128], and notions of theoretical rigor (of algorithmic and mathematical analysis) and empirical rigor (of statistical and experimental approaches). Theoretical rigor typically seeks precise problem formulation using well-defined mathematical notation, accompanied by results (e.g., propositions, theorems, lemmas) with correct proofs—e.g., a clear and correct sequence of mathematically derived steps that support the stated result. Empirical rigor, in turn, seeks comparison of a proposed algorithm with a sufficient number of alternative—often competing—approaches, ablation studies, and some form of statistical analysis (e.g., power analysis, significance tests, or simply error bars). Renewed calls for methodological rigor have often been motivated by reproducibility concerns [92, 49, 74], with checklists and documentation practices proposed as a way to support reproducibility and replicability by standardizing methodological choices, recording them, and making them explicit.
@@
Mechanisms to promote methodological rigor Methodological choices are shaped by considerations about how to operationalize what we know—e.g., the background knowledge—to substantiate existing knowledge or produce new knowledge or artifacts. As methodological concerns have been central to discussions about rigor in AI, here we only focus on foregrounding mechanisms for aspects of methodological rigor we believe deserve added attention, including construct validity [143, 26, 71] and the need for methodological standards, particularly for high-risk domains [149]. For more comprehensive discussions of methodological rigor we direct the reader to [e.g., 62, 129, 128, 143, 87, 74].
@@
Construct validity:  Many problems in AI, such as assessments of systems and phenomena, are concerned with measurement [71, 143]. Even when there is conceptual clarity, for these problems ensuring construct validity—i.e., that measurement instruments (e.g., benchmark metrics) appropriately capture the construct of interest (e.g., reasoning, understanding, values)—is foundational to meaningful measurement and thus methodological rigor. A growing body of work has proposed frameworks and best practices for assessing the validity of measurements [e.g., 106, 71, 140, 89] and illustrated that existing measurement instruments exhibit a range of concerns that threaten their ability to measure what they purport to measure [26, 59, 102]. For example, Northcutt et al. [102] show widespread label errors in benchmark test datasets which can “destabilize ML benchmarks,” thereby “lead[ing] practitioners to incorrect conclusions about which models actually perform best in the real world.”
@@
Compliance with methodological standards:  Establishing methodological standards often involves extensive, community-wide debates about what methods are appropriate and when. Petzschner [109] notes how the failure of ML models intended for medical settings “to generalize to data from new, unseen clinical trials […] highlight[s] the necessity for more stringent methodological standards,” particularly for high-risk settings [149]. This has precipitated calls for developing standards in health datasets in AI [9]. The standardization of information retrieval evaluation practices via NIST’s Text Retrieval Conference (TREC) was fundamental in revitalizing the research community and impacting the development of web search engines [118]. However, even though established standards can help a research community promote more rigorous debates about methodological choices, they may not by themselves ensure that methodological choices are explicitly reported or reflected on; Geiger et al. [53] hypothesize “that in fields with widely-established and shared methodological standards, researchers could have far higher rates of adherence to methodological best practices […] but have lower rates of reporting that they actually followed those practices.” By the same token, not making a choice of methods and presenting a kitchen sink (of metrics, methods) [155] undermines critical engagement with why the methods are appropriate. Compliance with standards should include explicit reflections on methodological choices and their application, including aspects related to constraints that researchers and practitioners had to navigate, such as access to participants, computing, or other resources [107].

### v2

@@ 2.2 Normative rigor
Because of differences in underlying evidentiary standards, background assumptions, goals, ideals, and theoretical frameworks, different disciplines and communities have different ways of determining what methods to use, what types of evidence are needed or sufficient, and why the resulting work or resulting knowledge matters [brister2017epistemological, carter2007justifying]. The assumptions and theories that underpin our work are not derived “out of thin air, but as a function of our experiences with the world in general and specifically with our colleagues, through dialogue, and the literature” [guest2024makes], all of which can shape the type, direction, and quality of research [howe2020stick, danziger1985methodological, george1973conceptualization]. Disciplinary and community norms are intertwined with a researcher’s personal norms and drivers [davies2002qualitative], e.g., what catches their interest or advances their goals. Making explicit how this mix of influences shapes the AI community’s work can help others understand them and enable debates on their appropriateness.
@@
Consider the emerging research on developing AI personas as tools to simulate users and human study participants [e.g., amin2025generative, lazik2025impostor, ge2024scaling, olteanu2025ai]. While this research reflects common norms and beliefs in AI communities around scalability and efficiency [e.g., birhane2022values, agnew2024illusion], by aiming to replace human participants, such tools may fail to align or may even conflict with foundational values and norms around representation, participation, inclusion, and understanding [agnew2024illusion, wang2025simulate] that underpin many types of work this research seeks to support, such as user experience or social science studies with human participants. Such value conflicts “cannot be alleviated with better training or improved model performance alone” [agnew2024illusion].
@@
Common beliefs and expectations in the AI community—such as the goal of producing generalizable findings that are often abstracted away from any use context [birhane2022values, hutchinson2022evaluation]—have led to a reduction of problems and use scenarios “to a common set of representations or affordances” [birhane2022values], and thus (even if implicitly) to de-valuing how context shapes datasets and tools. Reliance on undisclosed normative assumptions about problem statements, artifacts, or constructs that are contested or value-laden, yet treated as if they are neutral or generally applicable, can however be particularly misleading when a “field employs the language of procedural adherence to project a sense of certainty, objectivity, and stability” [green2018myth]. While dominant values in AI work like performance, generalization, and efficiency [birhane2022values] may encourage extensive evaluations across as many metrics, benchmarks, and baselines as possible [zhou2022deconstructing, birhane2022values], these evaluations tell us little about when, whether, or which improvements are necessary, desirable, or translate to meaningful benefits when deployed. Such expectations and norms—along resource-related considerations and constraints such as about costs and strict timeliness [zaenen2006last, zhou2022deconstructing, harvey2025understanding]—nonetheless influence what type of work gets prioritized or rewarded [fahimi2024articulation].
@@
Mechanisms to promote normative rigor All research is shaped by normative considerations. Normative rigor asks us to make these considerations explicit—such as via ethical and positionality statements [olteanu2023responsible] or other disclosure practices—and can include aspects related to expectations around the significance and impact of research [e.g., nowell2019reviewer, guest2024makes] or what ethical norms to follow [e.g., olteanu2019social].
@@
Research significance:  Central to normative considerations are expectations about how our work and its outcomes should intervene in the world, or why the work matters and is appropriate. The goal of research is often to “generate knowledge that will have positive practical impact” [ashurst2022disentangling] or that advances the field [nowell2019reviewer], with norms around what work is worthwhile driven by desires to reward such work [west2012rigor, burtoneditor]. There is however also an increasing recognition that the appraisal of the work’s quality and significance should also include the work’s potential for harm and not only for benefits [guest2024makes, boyarskaya2020overcoming, ashurst2022disentangling]. Such appraisal typically rests on claims about the possible impacts from either the work’s process or its outcomes. This echoes the concept of consequential validity from social science measurement scholarship [messick1987validity, wallach2025position], which asks us when determining a measurement instrument’s (in)validity to consider the consequential basis both of its use and of possible inferences (along with actions those inferences may entail) [messick1987validity, iliescu2021consequential]. Reckoning with the impact of our work thus promotes not only discussions about whether current norms around what constitutes good work are appropriate, but also promotes methodological (§2.4) and interpretative rigor (§2.6).
@@
Positionality and ethical statements:  Researchers’ personal, disciplinary, and institutional backgrounds, their lived experiences, and their goals motivate and shape how they approach their work. Positionality statements are a mechanism meant to make such considerations explicit in order to help others contextualize the research and research outcomes [olteanu2023responsible, liang2021reflexivity]. Positionality, however, does not only encompass aspects related to researchers’ beliefs and values, but also those related to what knowledge they draw on, how they know what they know, and how they make methodological choices [berkovic2023researcher]. Thus, it can also aid or compromise epistemic (§2.1) and methodological rigor (§2.4). Ethical statements can further complement positionality statements by foregrounding the ethical concerns researchers grappled with or mitigated before or while conducting the work [olteanu2023responsible, ashurst2022ai]. Yet this practice of disclosing whether such concerns were considered and how they shaped any methodological choices (if at all) remains inconsistently adopted across AI communities.
@@ 2.3 Conceptual rigor
Assume a researcher wishes to evaluate whether a model “hallucinates.” In AI research, the construct of “hallucination” has, however, been used to refer to several distinct types of system behaviors [MalekyEtAlHallucinations], including cases of generating content which is nonsensical, which contains factual errors, which is not in the input data, which is not in the training data, or a mix of these behaviors. Further, all these different understandings of what it means for a model to “hallucinate” are markedly different from the more common use of the term that requires an ability to perceive, feel, or have sensory experiences [e.g., halluciations-meaning]; and the term can thus carry meanings incompatible with AI systems. To understand what a researcher’s evaluation of whether the model “hallucinates” means, we thus need to know which conceptualization they use, and if that conceptualization is sensible.
@@
While contested constructs—those that have competing or even conflicting definitions, like “hallucination,” “value alignment,” “AGI,” or “human-likeness”—are increasingly common in AI research and practice, their definitions often remain elusive, ambiguous, or poorly specified [e.g. blodgett2021stereotyping, wallach2025position, blili2024unsocial, blili2025stop]. Without clarity about what specifically we are analyzing, measuring, or striving for, it can be hard to assess progress or make any useful or reliable claims. Work can also rely on constructs that are inconsistent with any theoretical tradition, such as treating identity categories as fixed and objective rather than continuous and constructed [lu2022subverting] or more generally failing to recognize that some constructs are innately fluid, non-deterministic, and fuzzy [birhane2021impossibility]. Such a lack of clarity can hinder replicability and reproducibility or may facilitate speculative post-factum interpretations, yielding possibly unsound and unfounded claims, or a conflation of proxy measurements (which may or may not measure any version of the underlying construct) with the construct under analysis. Thus, a lack of conceptual clarity can also undermining epistemic (§2.1), methodological (§2.4), and interpretative rigor (§2.6).
@@
Mechanisms to promote conceptual rigor Conceptual rigor requires attention to conceptual clarity—which construct we are after and how it is defined; appropriate conceptual systematization—the process by which the definition is made specific; and terminological rigor—that the terms used to refer to a construct do not harbor meanings that can lead to a misinterpretation of what the construct is.
@@
Conceptual clarity:  A growing number of objects in AI research are ambiguous or poorly specified, or else are objects for which we lack consensus about what they are or what they are for. The examination by saphra2024mechanistic of what is meant by “mechanistic interpretability” is instructive: not only does the term have multiple competing meanings, but those meanings also reflect distinct disciplinary orientations and epistemic origins; a lack of clarity about which meaning is under use can obfuscate not only what the work does, but also why and how the work is done. Similar critiques have been made about the lack of conceptual clarity about what unlearning [cooper2024unlearning], bias [blodgett-etal-2020-language], model collapse [schaeffer2025modelcollapse], interpretability [lipton2018mythos], or generalization [lipton2019troubling] mean. While we see growing concerns about the lack of conceptual clarity surrounding many aspects of AI research, from how desired capabilities are described to what metrics to optimize for [jacobs2021measurement, blodgett2021stereotyping, herrmann2024position, wallach2025position, saphra2024mechanistic], these remain largely overlooked in discussions about research integrity and quality in AI.
@@
Conceptual systematization:  In practice, many constructs involve a “broad constellation of meanings and understandings” [adcock2001measurement], and working with them requires making choices about which meanings to use, “narrowing [them] into an explicit definition” [wallach2025position]. The process of conceptual systematization asks researchers to engage not only with a high-level construct in the abstract, but to grapple more concretely with what it means in the context of the work they are conducting and how it relates to empirical observations or other constructs. Conceptual systematization is a prerequisite for rigorous measurement, (computational) specification, empirical analysis, and theory development [e.g., wallach2025position, porada2024challenges, adcock2001measurement, herrmann2024position], and is thus a prerequisite for methodological rigor (§2.4).
@@
Terminological rigor:  Conceptual rigor depends on terminological choices and what those choices communicate. Many terms in AI often carry over meanings from the human realm or other disciplinary contexts that are incompatible with AI systems, and can mislead [chen2025portraying, lipton2019troubling, blili2025stop, rehak2021language], suggest “unproven connotations,” or lead to “collisions with other definitions, or conflation with other related but distinct concepts” [lipton2019troubling]. blili2025stop note that “when researchers equate human faculties with model proxies […] [t]his rhetorical move is enabled by using colloquial terms like ‘imagination’ without considering whether it corresponds to the human faculty,” leading to inflated claims. Clear, precise language helps “dispel speculative, scientifically unsupported portrayals of [AI] systems, and support more factual descriptions of them” [cheng2024one], and thus clear scientific communication.
@@ 2.4 Methodological rigor
Rigor is often “conceptualized as the appropriate execution of [methods]” [nowell2019reviewer], with discussions about rigor in AI centering around methodological considerations, from data collection and analysis, to model training and tuning, to experimental practices [e.g., herrmann2024position, sculley2018winner, sculley2025position], and notions of theoretical rigor (of algorithmic and mathematical analysis) and empirical rigor (of statistical and experimental approaches). Theoretical rigor typically seeks precise problem formulation using well-defined mathematical notation, accompanied by results (e.g., propositions, theorems, lemmas) with correct proofs—e.g., a clear and correct sequence of mathematically derived steps that support the stated result. Empirical rigor, in turn, seeks comparison of a proposed algorithm with a sufficient number of alternative—often competing—approaches, ablation studies, and some form of statistical analysis (e.g., power analysis, significance tests, or simply error bars). Renewed calls for methodological rigor have often been motivated by reproducibility concerns [magnusson2023reproducibility, forde2019scientific, kapoor2024reforms], with checklists and documentation practices proposed as a way to support reproducibility and replicability by standardizing methodological choices, recording them, and making them explicit.
@@
Mechanisms to promote methodological rigor Methodological choices are shaped by considerations about how to operationalize what we know—e.g., the background knowledge—to substantiate existing knowledge or produce new knowledge or artifacts. As methodological concerns have been central to discussions about rigor in AI, here we only focus on foregrounding mechanisms for aspects of methodological rigor we believe deserve added attention, including construct validity [wallach2025position, blodgett2021stereotyping, jacobs2021measurement] and the need for methodological standards, particularly for high-risk domains [wilhelm2025benefits]. For more comprehensive discussions of methodological rigor we direct the reader to [e.g., herrmann2024position, sculley2018winner, sculley2025position, wallach2025position, lipton2019troubling, kapoor2024reforms].
@@
Construct validity:  Many problems in AI research, such as assessments of systems and phenomena, are concerned with measurement [jacobs2021measurement, wallach2025position, mitchell2023measuringdata]. Even when there is conceptual clarity, ensuring construct validity—i.e., that measurement instruments (e.g., benchmark metrics) appropriately capture the construct of interest (e.g., reasoning, understanding, values)—is foundational to meaningful measurement and thus methodological rigor. A growing body of work has proposed frameworks and best practices for assessing the validity of measurements [e.g., olteanu2019social, jacobs2021measurement, wagner2021measuring, liu-etal-2024-ecbd] and illustrated that existing measurement instruments exhibit a range of concerns that threaten their ability to measure what they purport to measure [blodgett2021stereotyping, gururangan-etal-2018-annotation, northcutt2021pervasive]. For example, northcutt2021pervasive show widespread label errors in benchmark test datasets which can “destabilize ML benchmarks,” thereby “lead[ing] practitioners to incorrect conclusions about which models actually perform best in the real world.”
@@
Compliance with methodological standards:  Establishing methodological standards often involves extensive, community-wide debates about what methods are appropriate and when. petzschner2024practical notes how the failure of ML models intended for medical settings “to generalize to data from new, unseen clinical trials […] highlight[s] the necessity for more stringent methodological standards,” particularly for high-risk settings [wilhelm2025benefits]. This has precipitated calls for developing standards in health datasets in AI applications [arora2023value]. The standardization of information retrieval evaluation practices via NIST’s Text Retrieval Conference (TREC) was fundamental in revitalizing the research community and impacted the development of web search engines [trec-economic-impact-report]. However, even though established standards can help a research community promote more rigorous debates about methodological choices, they may not by themselves ensure that those choices are explicitly reported or reflected on; geiger2021garbage hypothesize “that in fields with widely-established and shared methodological standards, researchers could have far higher rates of adherence to methodological best practices […] but have lower rates of reporting that they actually followed those practices.” By the same token, not making a choice of methods and presenting a kitchen sink (of metrics, methods) [zhou2022deconstructing] undermines critical engagement with why the methods are appropriate. Compliance with standards should include explicit reflections on methodological choices and their application, including aspects related to constraints that researchers and practitioners had to navigate, such as access to participants, computing, or other resources [olteanu2023responsible].